# “Ask Me Anything”: How Comcast Uses LLMs to Assist Agents in Real Time

### v1

@@ Abstract.
Customer service is how companies interface with their customers. It can contribute heavily towards the overall customer satisfaction. However, high-quality service can become expensive, creating an incentive to make it as cost efficient as possible and prompting most companies to utilize AI-powered assistants, or ”chat bots”. On the other hand, human-to-human interaction is still desired by customers, especially when it comes to complex scenarios such as disputes and sensitive topics like bill payment. [^1] https://www.businessinsider.in/tech/enterprise/news/why-customers-still-prefer-speaking-to-a-human-not-chatbox/articleshow/91955173.cms
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3661345isbn: 979-8-4007-0431-4/24/07ccs: Information systems Retrieval models and rankingccs: Information systems Language models

### v2

@@ Abstract.
Customer service is how companies interface with their customers. It can contribute heavily towards the overall customer satisfaction. However, high-quality service can become expensive, creating an incentive to make it as cost efficient as possible and prompting most companies to utilize AI-powered assistants, or ”chat bots”. On the other hand, human-to-human interaction is still desired by customers, especially when it comes to complex scenarios such as disputes and sensitive topics like bill payment. [^1] https://bit.ly/3yaNO9t
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAdoi: 10.1145/3626772.3661345isbn: 979-8-4007-0431-4/24/07ccs: Information systems Retrieval models and rankingccs: Information systems Language models